2 kg of mass is placed on x axis 2m away from the origin, the particle is released from rest to fall under the effect of gravity. Calculate the angular momentum and torque


1
Expert's answer
2021-04-08T09:02:13-0400

At "t=0" the angular momentum is "L=0".


Torque "\tau=r\cdot F=r\cdot mg=2\cdot2\cdot9.81=39.24\ (N\cdot m)" .